If angle B measures forty-seven degrees, what are the measures of angle A and angle C, given that angle A and angle B are supplementary, and angle B and angle C are complimentary?
Answer by stanbon(75887)

If angle B measures forty-seven degrees, what are the measures of angle A and angle C, given that angle A and angle B are supplementary, and angle B and angle C are complimentary?
------
Equations:
B = 180-A
B = 90-C
B = 47 degrees
---------------------
A = 180-B = 180-47 = 133 degrees
C = 90-B = 90-47 = 43 degrees
==================================
cheers,
Stan H.